# Diff-Reg v1: Diffusion Matching Model for Registration Problem

### v3

@@ III-E The Lightweight Denoising Module $g_{\theta}$
This section outlines the architecture of the lightweight denoising module $g_{\theta}$. In 3D registration tasks, we take the downsampled superpoints $\hat{P}$ and $\hat{Q}$ along with the points features $F^{\hat{P}}$ and $F^{\hat{Q}}$ as the inputs of $g_{\theta}$ in each reverse sampling step. Similarly, in 2D3D registration tasks, the inputs of $g_{\theta}$ are downsampled superpixels $\hat{X}$ and $\hat{Q}$ with associated inputs $\{F^{\hat{X}},D^{\hat{X}},F^{\hat{X}}_{dino},F^{\hat{Q}}\}$. When inference, $g_{\theta}$ inputs a noised matching matrix $E^{t}$ and outputs a predicted target matching matrix $\hat{E}_{0}$.
We define the denoising module $g_{\theta}$ by sequentially stacking five components as a differentiable layer: SinkHorn Projection, Weighted SVD, Warping Function, Denoising Transformer, and Matching function. The Weighted SVD operation can be viewed as a differentiable projection that arises from the matching matrix obtained from the first Sinkhorn Projection. Alternatively, in the context of a 2D-3D registration task, this Weighted SVD operation can be substituted with a differentiable PnP (Perspective-n-Point) solver[55].
@@
Sinkhorn Projection: $\mathbf{f}_{\text{sinkhorn}}(\cdot)$. To constrain the matching matrix $E^{t}$ within the doubly stochastic matrices manifolds, we utilize the SinkHorn[50] iterations to project $E^{t}$. We treat this operation as a key role in our framework rather than a post-processing in other methods.
@@
Weighted SVD: $\mathbf{soft\_procrustes}(\cdot,\cdot,\cdot)$. Given top-k confident correspondences $K$, we utilize the weighted SVD algorithm[56] (differentiable) to compute the transformation $R,t$ in a closed form:
@@
$\begin{split}H&=\sum\limits_{(i,j)\in C}\tilde{E}(i,j)\hat{p}_{i}\hat{q}_{j}^{T},\ H=U{\Sigma}V^{T}\\ \mathbf{R}&=Udiag(1,1,det(UV^{T}))V,\\ \mathbf{t}&=\frac{1}{|K|}\left({\sum\limits_{(i,\cdot)\in K}\ \hat{p}_{i}}-\mathbf{R}{\sum\limits_{(\cdot,j)\in K}}\hat{q}_{j}\right)\end{split}$ (6)
@@
where $(\hat{p}_{i},\hat{q}_{j})$ is a superpoint correspondences. In the 2D3D registration task, we first project a superpixel depth map $D^{\hat{X}}$ to a point cloud $P^{\hat{X}}_{D}$ using camera intrinsic. Then we utilize this SVD decomposition to compute $\mathbf{R},\mathbf{t}$ with $P^{\hat{X}}_{D}$ and $\hat{Q}$ as inputs.
@@
Warping Function: $\mathbf{warping}(\cdot,\cdot,\cdot)$. After obtaining transformation $\mathbf{R},\mathbf{t}$, the rigid warping of source point clouds is computed by $W(\hat{p}_{i})=\mathbf{R}\hat{p}_{i}+\mathbf{t}$. In this paper, we use rigid warping for rigid and deformable registration cases to demonstrate our design. As for deformable registration, with the denoised correspondences, we can compute the flow fields for all points in $\hat{P}$ by performing nearest neighbor interpolation with the predicted inlier correspondences as anchors. For the deformation warping, we actually can build a deformation graph[57, 58] to conduct deformable warping[30, 59] for the source point cloud $\hat{P}$. We leave this deformable warping integration in future work.
@@
Denoising Transformer: $f_{\theta}(\cdot,\cdot,\cdot,\cdot,\cdot,\cdot)$. We observed empirically that a simple noise model does not hurt performance. Thus, we exploit a lightweight Transformer[54] as our denoising network. Specifically, we utilize a 6-layer inter-leaved attention layers transformer $f_{\theta}$ for denoising feature embedding.
@@
Attention layer.. In the 3D registration task, following [8], the $q,k,v$ in the self-attention, are computed as:
$q_{i}=\Theta(p_{i})\mathbf{W}_{q}f^{\hat{p}_{i}},\ k_{j}=\Theta(p_{j})\mathbf{W}_{k}f^{\hat{p}_{j}},\ v_{j}=\mathbf{W}_{v}f^{\hat{p}_{j}},$ (7)
$f^{\hat{p}_{i}}=f^{\hat{p}_{i}}+\text{MLP}(\text{cat}[f^{\hat{p}_{i}},\Sigma_{j}\alpha_{ij}v_{j}]),$
@@
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in{\mathbb{R}}^{d{\times}d}$ are the attention weights, ${\alpha}_{ij}=softmax(q_{i}k_{j}^{T}/{\sqrt{d}})$, and $\Theta(\cdot)$ is a rotationary position encoding [8]. MLP()($\cdot$)( ⋅ ) is a 3-layer fully connected network, and $cat[{\cdot},\cdot]$ is the concatenating operator. The cross attention layer is the standard form that $q$ and $k,v$ are computed by source and target point clouds, respectively. Other operations are the same with self-attention. In the 2D3D registration task, we take image inputs $\{\hat{X},F^{\hat{X}},F^{\hat{X}}_{dino}\}$ and point cloud input $\{\hat{Q},F^{\hat{Q}}\}$ to compute $q,k,v$ by utilizing standard attention layers[54]. We also take Fourier embedding function[60] to embed superpixels $\hat{X}$ and superpoints $\hat{Q}$ as positional encoding.
@@
Matching function: $\mathbf{matching\_logits}(\cdot,\cdot,\cdot,\cdot)$. We compute matching “logits”between $\hat{P}$ and $\hat{Q}$ by features $F^{\hat{P}}$ (or $F^{\hat{X}}$) and $F^{\hat{Q}}$: $\tilde{E}(i,j)=\frac{1}{\sqrt{d}}\left<f^{\hat{p}_{i}},f^{\hat{q}_{j}}\right>$.
@@
For the sake of clarity, we provide pseudo-code in Algorithm.1 to describe the logic of our entire denoising module $g_{\theta}$ for the 3D Registration task. The definition of $g_{\theta}$ for the 2D3D registration task is similar, which can be found in section (IV-B), Algorithm.4, and Algorithm.5.
@@
IV Experiments
In this section, we conduct experiments on 3D registration and 2D3D registration tasks to prove the effectiveness of our method.
Figure 3: Overview of our diffusion matching model in 3D registration task. The top half is the single-pass prediction head with the feature backbone, while the bottom is our denoising module.
Algorithm 2 Training Diff-Reg in 3D Registration Task
1:Point clouds $\hat{\mathbf{P}}$, $\hat{\mathbf{Q}}\in\mathbb{R}^{3}$ and associated point features $\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}}$.
2:while not converged do
3:     Sample $\mathbf{E}^{0}\sim q(\mathbf{E}^{0})$
4:     $N\times M\leftarrow\mathbf{E}^{0}.\text{shape}$
5:     Sample $t\sim\text{Uniform}(1,...,T)$
6:     $\epsilon\sim\mathcal{N}(0,1)^{N\times M}$
7:     if Rigid then
8:         $\mathbf{E}^{t}\leftarrow\sqrt{\bar{\alpha}_{t}}\mathbf{E}^{0}+\sqrt{1-\bar{\alpha}_{t}}f_{\epsilon}(\epsilon),\quad\tilde{\mathbf{E}}^{t}=\mathbf{E}^{t}-\rm{Min}(\mathbf{E}^{t})$
9:     else
10:         $\mathbf{E}^{t}=\sqrt{\bar{\alpha_{t}}}\mathbf{E}^{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon_{0},\quad\tilde{\mathbf{E}}^{t}=\rm{Sigmoid}(\mathbf{E}^{t})$
11:     end if
12:     $\hat{\mathbf{E}}_{0}\leftarrow g_{\theta}(\tilde{\mathbf{E}}^{t},\hat{\mathbf{P}},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}})$
13:     Optimize $L_{t}=\rm{Focal\_loss}(\hat{\mathbf{E}}_{0},\mathbf{E}^{0})$
14:end while
Algorithm 3 Sampling by Diff-Reg in 3D Registration Task
1:Initial matching matrix $E^{T}$ from backbone or white noise; Point clouds $\hat{\mathbf{P}},\hat{\mathbf{Q}}\in\mathbb{R}^{3}$ and associated point features $\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}}$.
2:Target matching matrix $\tilde{\mathbf{E}}^{0}$.
3:$N\times M\leftarrow\mathbf{E}^{T}.\text{shape}$
4:for $t=T,...,1$ do
5:     ${\mathbf{z}_{t}}\sim N(0,1)^{N\times M}$ if $t>1$ else ${\mathbf{z}_{t}}\leftarrow\mathbf{0}^{N\times M}$
6:     if Rigid then
7:         $\tilde{\mathbf{E}}^{t}=\mathbf{E}^{t}-\rm{Min}(\mathbf{E}^{t})$
8:     end if
9:     $\hat{\mathbf{E}}_{0}\leftarrow g_{\theta}(\hat{\mathbf{E}}^{t},\hat{\mathbf{P}},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{P}}},\mathbf{F}^{\hat{\mathbf{Q}}})$
10:     $\epsilon_{t}\leftarrow\frac{\hat{\mathbf{E}}_{0}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\mathbf{E}}^{t}$
11:     $\sigma_{t}\leftarrow\sqrt{\frac{(1-\alpha_{t-1})}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-1}}}$
12:     $\hat{\mathbf{E}}^{t-1}\leftarrow\sqrt{\alpha_{t-1}}\hat{\mathbf{E}}_{0}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon_{t}+\sigma_{t}\mathbf{z}_{t}$
13:end for
14:if Rigid then
15:     $\tilde{\mathbf{E}}^{0}=\tilde{\mathbf{E}}^{0}-\rm{Min}(\tilde{\mathbf{E}}^{0}),\quad\tilde{\mathbf{E}}^{0}=\mathbf{f}_{\rm{sinkhorn}}(\tilde{\mathbf{E}}^{0})$
16:else
17:     $\tilde{\mathbf{E}}^{0}=\rm{Sigmoid}(\tilde{\mathbf{E}}^{0})$
18:end if
IV-A 3d Registration Task
IV-A1 Implementation Details.
The training of our 3D registration network architecture is shown in Fig. 3. We utilize the joint training strategy for the single-pass prediction head and our denoising module. Our framework utilize a KPConv[14] branch to give the downsampled the superpoints $\hat{P}$ and $\hat{Q}$ and the asccociated features $F^{\hat{P}}$ and $F^{\hat{Q}}$. Subsequently, we employ a repositioning transformer [8] to provide a single-pass prediction of the matching matrix and the resulting transformation $R,t$ both of which are supervised by the Matching loss $L_{M}$ and Warping loss $L_{W}$ introduced in [8]. We utilize the focal loss $L_{simple}$ (Eqn.5) to guide the training of the denoising module $g_{\theta}$. The total loss function is defined as $L=L_{M}+L_{W}+L_{simple}$. During inference, we use the feature backbone to generate the superpoint features $F^{\hat{P}}$ and $F^{\hat{Q}}$ and treat them as constant frozen features in the reverse denoising sampling process. The dimension $d$ of superpoint features $F^{\hat{P}}$ and $F^{\hat{Q}}$ is set as $d=528$. The training and reverse sampling of denoising module $g_{\theta}$ are in Algorithm. 2 and Algorithm. 3.
Our proposed model is trained and evaluated using PyTorch on an NVIDIA RTX 3090 GPU. We train the model for approximately 30 epochs on the 3DMatch and 4DMatch datasets with a batch size of 2. We adopt the training/validation/test split strategy from [11, 8] for 3DMatch and 4DMatch datasets. The total diffusion steps during training are set to 1000.
IV-A2 Non-Rigid 4DMatch/4DLoMatch Benchmark.
Datasets.
@@ IV Experiments
4DMatch/4DLoMatch[8] is an benchmark generated by the animation sequences from DeformingThings4D[61]. We follow the dataset split provided in [8], which has a wide range of overlap ratio, that $45\%$-$92\%$ in 4DMatch and $15\%$-$45\%$ in 4DLoMatch.
Metrics.
@@
Following Lepard [8], we utilize two evaluation metrics to assess the quality of predicted matches. (1) Inlier ratio (IR): This metric represents the correct fraction in the correspondences prediction $\mathcal{K}_{pred}$. (2) Non-rigid Feature Matching Recall (NFMR): This metric calculates the fraction of ground truth correspondences $(u,v)\in\mathcal{K}_{gt}$ that can be successfully recovered by using the predicted correspondences $\mathcal{K}_{pred}$ as anchors. The NFMR metric provides a better characterization of the global rationality of overall body deformation, directly indicating whether the anchor $\mathcal{K}_{pred}$ effectively captures the body movements.
Specifically, IR measure denotes the correct fraction in the correspondences prediction $\mathcal{K}_{pred}$:
$IR=\frac{1}{|\mathcal{K}_{pred}|}\Sigma_{(p,q)\in\mathcal{K}_{pred}}[||W_{gt}(p)-q||_{2}<\sigma]$ (8)
where $||\cdot||_{2}$ is the Euclidean norm, $W_{gt}(\cdot)$ is the ground truth warping function, []delimited-[][$\cdot$][ ⋅ ] is the Inverse bracket, and $\sigma=0.04m$. The NFMR metric is to compute the fraction of the ground truth correspondences $(u,v)\in\mathcal{K}_{gt}$ that can be successfully recovered from the predicted correspondences $\mathcal{K}_{pred}$. First, we construct the predicted correspondences $\mathcal{A}=\{p|(p,q)\in\mathcal{K}_{pred}\}$ and the associated sparse 3D flow fields =$\mathcal{F}=\{q-p|(p,q)\in\mathcal{K}_{pred}\}$. Then, for any source point $u$ in $\mathcal{K}_{gt}$, we can recover the flow field for $u$ by inverse distance interpolation:
$\Gamma(u,\mathcal{A},\mathcal{F})=\Sigma_{\mathcal{A}_{i}\in\mathcal{N}(u,\mathcal{A})}\frac{\mathcal{F}_{i}||u-\mathcal{A}_{i}||^{-1}_{2}}{\Sigma_{\mathcal{A}_{i}\in\mathcal{N}(u,\mathcal{A})}||u-\mathcal{A}_{i}||_{2}^{-1}}$ (9)
where $\mathcal{N}(\cdot,\cdot)$ is k-nearest neighbors search with k = 3. After that, we define the $NFMR$ to measure the fraction of ground truth matches that we recovered from $\mathcal{K}_{pred}$:
$NFMR=\frac{1}{|\mathcal{K}_{gt}|}[||\Gamma(u,\mathcal{A},\mathcal{F})-(v-u)||_{2}<\sigma]$ (10)
Quantitative results.
We compare our method with two categories of state-of-the-art methods. The first category includes Scene Flow Methods such as PWC[62], FLOT[63], and NSFP[64]. The second category encapsulates Feature Matching-Based Methods, namely D3Feat[3], Predator[11], Lepard[8], GeoTR[6], and RoITr[2]. As illustrated in Table I, our method demonstrates significant improvements compared to the single-pass baselines. “Diff-Reg(Backbone)” refers to the single-pass prediction head (i.e., reposition transformer in Lepard [8]), while “Diff-Reg(Diffusion/1)” and “Diff-Reg(Diffusion/20)” utilize our denoising module $g_{\theta}$ with one single step and 20 steps of reverse denoising sampling. For both NFMR and IR metrics, “Diff-Reg(Diffusion/20)” achieves the best performance. The improvement in NFMR of “Diff-Reg(Backbone)” compared to the baselines indicates that our diffused training samples in the matching matrix space enhance the feature backbone’s representation, enabling the capture of crucial salient correspondences that are helpful for consistent global deformation. The significant enhancement of “Diff-Reg(Diffusion/20)” over “Diff-Reg(Diffusion/1)” demonstrates that the reverse denoising sampling process indeed searches for a better solution guided by the learned posterior distribution. To validate that the predicted correspondences indeed improve deformable registration, we conducted experiments using the state-of-the-art registration method GraphSCNet [59]. As indicated in Table II, our predicted correspondences are beneficial for deformable registration, particularly in the more challenging 4DLoMatch benchmark.
@@
TABLE I: Quantitative results on the 4DMatch and 4DLoMatch benchmarks. The best results are highlighted in bold, and the second-best results are underlined.
@@
Diff-Reg(Diffusion/1) 85.23 83.85 73.19 65.26
Diff-Reg(Diffusion/20) 88.40 86.41 76.23 67.80
TABLE II: Non-rigid registration results of 4DMatch/4DLoMatch. Given predicted correspondences, we utilize the non-rigid registration method GraphSCNet[59] to conduct the deformable registration. We retrain RoITr using the authors’ code.
@@
Figure 4: The qualitative results of non-rigid registration in the 4DMatch/4DLoMatch benchmark. The top two lines are from 4DMatch, while the bottom three are from 4DLoMatch. The blue and yellow colors denote the source and target point cloud, respectively. The green and red lines indicate whether the threshold accepts the predicted deformable flow from the source points. The deformable registration is built by GraphSCNet[4]. Zoom in for details.
TABLE III: Quantitative results on the 3DMatch and 3DLoMatch benchmarks. The best results are highlighted in bold, and the second-best results are underlined.
Method 3DMatch 3DLoMatch
[…]
FCGF ICCV2019 [65] 95.20 56.90 88.20 60.90 21.40 45.80
D3Feat CVPR2020 [3] 95.80 39.00 85.80 69.30 13.20 40.20
Predator CVPR2021 [11] 96.70 58.00 91.80 78.60 26.70 62.40
Lepard CVPR2022 [8] 97.95 57.61 93.90 84.22 27.83 70.63
GeoTR CVPR2022 [6] 98.1 72.7 92.3 88.7 44.7 75.4
RoITr CVPR2023 [2] 98.0 82.6 91.9 89.6 54.3 74.8
PEAL-3D CVPR2023 [33] 98.5 73.3 94.2 87.6 49.0 79.0
Diff-Reg(diffusion/1) 96.28 30.92 94.8 69.6 9.6 73.3
Diff-Reg(diffusion/20) 96.28 30.92 95.0 69.6 9.6 73.8
Qualitative results.

### v4

@@ 3.4 The Lightweight Denoising Module $g_{\theta}$
This section outlines the architecture of the lightweight denoising module $g_{\theta}$. In 3D registration tasks, $g_{\theta}$ take the superpoints $\hat{\mathbf{P}}$, $\hat{\mathbf{Q}}$ with associated points features $\mathbf{F}^{\hat{\mathbf{P}}}$, $\mathbf{F}^{\hat{\mathbf{Q}}}$ as the inputs in the reverse sampling process. Similarly, in 2D-3D registration tasks, the inputs of $g_{\theta}$ are superpixels $\hat{\mathbf{X}}$ and superpoints $\hat{\mathbf{Q}}$ with associated features $\{\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{D}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino},\mathbf{F}^{\hat{\mathbf{Q}}}\}$. At inference time, $g_{\theta}$ inputs a noised matching matrix $\mathbf{E}^{t}$ and outputs a predicted target matching matrix $\hat{\mathbf{E}}_{0}$.
We define the denoising module $g_{\theta}$ by sequentially stacking five components as a differentiable layer:
@@
Sinkhorn Projection: $\mathbf{f}_{\text{sinkhorn}}(\cdot)$. To constrain the matching matrix $\mathbf{E}^{t}$ within the doubly stochastic matrices manifolds, we utilize the SinkHorn [12] iterations to project $\mathbf{E}^{t}$. We treat this operation as a key role in our framework rather than a post-processing in other methods.
@@
Weighted SVD: $\mathbf{soft\_procrustes}(\cdot,\cdot,\cdot)$. Given top-k confident correspondences $\kappa$, we utilize the weighted SVD algorithm [1] (differentiable) to compute the transformation $\mathbf{R},\mathbf{t}$ in a closed form:
@@
$\begin{split}\mathbf{H}&=\sum\limits_{(i,j)\in C}\tilde{\mathbf{E}}(i,j)\hat{\mathbf{p}}_{i}\hat{\mathbf{q}}_{j}^{\top},\ \mathbf{H}=\mathbf{U}{\mathbf{\Lambda}}\mathbf{V}^{\top},\\ \mathbf{R}&=\mathbf{U}{\rm diag}(1,1,\det(\mathbf{U}\mathbf{V}^{\top}))\mathbf{V},\\ \mathbf{t}&=\frac{1}{|\kappa|}\left({\sum\limits_{(i,\cdot)\in\kappa}\ \hat{\mathbf{p}}_{i}}-\mathbf{R}{\sum\limits_{(\cdot,j)\in\kappa}}\hat{\mathbf{q}}_{j}\right)\end{split}$ (6)
@@
where $(\hat{\mathbf{p}}_{i},\hat{\mathbf{q}}_{j})$ is a superpoint correspondences. In the 2D-3D registration task, we first project a superpixel depth map $\mathbf{D}^{\hat{\mathbf{X}}}$ to a point cloud $\mathbf{P}^{\hat{\mathbf{X}}}_{D}$ using camera intrinsic. Then we utilize this SVD decomposition to compute $\mathbf{R},\mathbf{t}$ with $\mathbf{P}^{\hat{\mathbf{X}}}_{D}$ and $\hat{\mathbf{Q}}$ as inputs.
The Weighted SVD operation can be viewed as a differentiable projection that arises from the matching matrix $\tilde{\mathbf{E}}$ obtained from the first Sinkhorn Projection. Alternatively, in the 2D-3D registration task, this Weighted SVD operation can be substituted with a differentiable PnP (Perspective-n-Point) [26].
@@
Warping Function: $\mathbf{warping}(\cdot,\cdot,\cdot)$. After obtaining transformation $\mathbf{R},\mathbf{t}$, the rigid warping of source point clouds is computed by $\mathbf{W}(\hat{\mathbf{p}}_{i})=\mathbf{R}\hat{\mathbf{p}}_{i}+\mathbf{t}$. In this paper, we use rigid warping for rigid and deformable registration cases to demonstrate our design. As for deformable registration, with the denoised correspondences, we can compute the flow fields for all points in $\hat{\mathbf{P}}$ by performing nearest neighbor interpolation with the predicted inlier correspondences as anchors.
@@
Denoising Transformer: $f_{\theta}(\cdot,\cdot,\cdot,\cdot,\cdot,\cdot)$. We observed empirically that a simple noise model does not hurt performance. Thus, we exploit a lightweight Transformer [47] as our denoising network. Specifically, we utilize a 6-layer inter-leaved attention layers transformer $f_{\theta}$ for denoising feature embedding. Worth noting that, in each denoising step, only coarse level source point cloud $\hat{\mathbf{P}}_{t}$ and its position encoding $\Theta(\hat{\mathbf{P}}_{t})$ (or target point cloud $\hat{\mathbf{Q}}$ in 2D-3D registration task) have their values changed according to the warping operation, while other input parameters remain fixed. This is the key to our fast sampling speed.
@@
Attention Layer in $f_{\theta}$: In the 3D registration task, following [28], the vectors $\mathbf{q},\mathbf{k},\mathbf{v}$ in the self-attention, are computed as:
$\mathbf{q}_{i}=\Theta(\mathbf{p}_{i})\mathbf{W}_{q}f^{\hat{\mathbf{p}}_{i}},\ \mathbf{k}_{j}=\Theta(\mathbf{p}_{j})\mathbf{W}_{k}f^{\hat{\mathbf{p}}_{j}},\ \mathbf{v}_{j}=\mathbf{W}_{v}f^{\hat{\mathbf{p}}_{j}},$ (7)
$f^{\hat{\mathbf{p}}_{i}}=f^{\hat{\mathbf{p}}_{i}}+\rm{MLP}(\text{cat}[\mathbf{q}_{i},\Sigma_{j}\alpha_{ij}\mathbf{v}_{j}]),$
@@
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in{\mathbb{R}}^{d{\times}d}$ are the attention weights, ${\alpha}_{ij}=softmax(\mathbf{q}_{i}\mathbf{k}_{j}^{\top}/{\sqrt{d}})$, and $\Theta(\cdot)$ is the relative rotationary position encoding [28]. MLP()($\cdot$)( ⋅ ) is a 3-layer fully connected network, and $cat[{\cdot},\cdot]$ is the concatenating operator. The cross attention layer is the standard form that $\mathbf{q}$ and $\mathbf{k},\mathbf{v}$ are computed by source and target point clouds, respectively.
In the 2D-3D registration task, we take image inputs $\{\hat{\mathbf{X}},\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino}\}$ and point cloud input $\{\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{Q}}}\}$ to compute $\mathbf{q},\mathbf{k},\mathbf{v}$ by utilizing standard attention layers [47]. We also take Fourier embedding function [35] to embed superpixels $\hat{\mathbf{X}}$ and superpoints $\hat{\mathbf{Q}}$ for positional encoding.
@@
Matching Function: $\mathbf{matching\_logits}(\cdot,\cdot,\cdot,\cdot)$. We compute matching “logits” between $\hat{\mathbf{P}}$ and $\hat{\mathbf{Q}}$ by features $\mathbf{F}^{\hat{\mathbf{P}}}$ (or $\mathbf{F}^{\hat{\mathbf{X}}}$) and $\mathbf{F}^{\hat{\mathbf{Q}}}$: $\tilde{\mathbf{E}}(i,j)=\frac{1}{\sqrt{d}}\left<f^{\hat{\mathbf{p}}_{i}},f^{\hat{\mathbf{q}}_{j}}\right>$.
@@
For the sake of clarity, we provide pseudo-code in Algorithm.1 to describe the logic of our entire denoising module $g_{\theta}$ for the 3D Registration task. A similar definition of $g_{\theta}$ for the 2D-3D registration task can be found in the appendix.
@@
4 Experiments
4.1 3D Non-Rigid Registration Task
@@ 4 Experiments
Datasets. 4DMatch/4DLoMatch [28] is an 3D non-rigid benchmark generated by the animation sequences from DeformingThings4D [30]. We follow the dataset split provided in [28], which has a wide range of overlap ratio, that $45\%$-$92\%$ in 4DMatch and $15\%$-$45\%$ in 4DLoMatch.
4.1.1 Implementation Details.
Our framework utilize a KPConv [45] backbone to produce the superpoints $\hat{\mathbf{P}}$ and $\hat{\mathbf{Q}}$ and the asccociated features $\mathbf{F}^{\hat{\mathbf{P}}}$ and $\mathbf{F}^{\hat{\mathbf{Q}}}$. The dimension $d$ of superpoint features $\mathbf{F}^{\hat{\mathbf{P}}}$ and $\mathbf{F}^{\hat{\mathbf{Q}}}$ is set as $d=432$. Subsequently, we employ a repositioning transformer [28] to provide a single-pass prediction of the matching matrix and the resulting transformation $[\mathbf{R},\mathbf{t}]$, both of which are supervised by the matching loss $L_{M}$ and warping loss $L_{W}$ introduced in [28]. We utilize a focal loss $L_{simple}$ (modified from Eqn.5) to guide the training of the denoising module $g_{\theta}$. The total loss function is defined as $L=L_{M}+L_{W}+L_{simple}$.
We train the model for 30 epochs on the 4DMatch dataset with a batch size of 2. We adopt the training/validation/test split strategy from Predator [19] and Lepard [28]. At inference time, we conduct 20 iterations in the reverse sampling process while the total diffusion steps during training are set to 1000.
@@
Metrics. Following Lepard [28], we utilize two evaluation metrics to assess the quality of predicted matches. (1) Inlier Ratio (IR): The correct fraction in the correspondences prediction $\mathcal{K}_{pred}$. (2) Non-rigid Feature Matching Recall (NFMR): The fraction of ground truth correspondences $(u,v)\in\mathcal{K}_{gt}$ that can be successfully recovered by using the predicted correspondences $\mathcal{K}_{pred}$ as anchors. The NFMR metric provides a better characterization of the global rationality of overall body deformation, directly indicating whether the anchor $\mathcal{K}_{pred}$ effectively captures the body movements.
@@
Table 1: Quantitative results on the 4DMatch and 4DLoMatch benchmarks. The best results are highlighted in bold, and the second-best results are underlined.
@@
Diff-Reg(steps=1) 85.23 83.85 73.19 65.26
Diff-Reg(steps=20) 88.40 86.41 76.23 67.80
Quantitative Results. We compare our method with two categories of state-of-the-art methods. The first category includes Scene Flow Methods such as PWC [55], FLOT [38], and NSFP [27]. The second category encapsulates Feature Matching-Based Methods, namely D3Feat [4], Predator [19], Lepard [28], GeoTR [40], and RoITr [66].
As illustrated in Table 1, our method demonstrates significant improvements compared to the single-pass baselines. “Diff-Reg(Backbone)” refers to the single-pass prediction head (i.e., reposition transformer in Lepard [28]), while “Diff-Reg(steps=1)” and “Diff-Reg(steps=20)” denotes our denoising module $g_{\theta}$ with one single step and 20 steps of reverse sampling. For both NFMR and IR metrics, “Diff-Reg(steps=20)” achieves the best performance. The improvement in NFMR of “Diff-Reg(Backbone)” compared to the baselines indicates that our diffused training samples in the matching matrix space enhance the feature backbone’s representation, enabling the capture of crucial salient correspondences that are helpful for consistent global deformation. The significant enhancement of “Diff-Reg(steps=20)” over “Diff-Reg(steps=1)” demonstrates that the reverse denoising sampling process indeed searches for a better solution guided by the learned posterior distribution.
To validate that the predicted correspondences indeed improve deformable registration, we conducted experiments using the state-of-the-art registration method GraphSCNet [41]. As indicated in Table 2, our predicted correspondences are beneficial for deformable registration, particularly in the more challenging 4DLoMatch benchmark.
Table 2: Non-rigid registration results of 4DMatch/4DLoMatch. Given predicted correspondences, we utilize the non-rigid registration method GraphSCNet [41] to conduct the deformable registration. We retrain RoITr using the authors’ code. The modified 4DMatch-F and 4DLoMatch-F datasets [29] exclude data involving near-rigid movements. The metrics (following [41, 29]) are 3D End Point Error (EPE), 3D Accuracy Strict (AccS) (<2.5cm or 5%), 3D Accuracy Relaxed (AccR)(<5cm or 5%), and Outlier Ratio (OR) ( >30%).
@@
Figure 2: The qualitative results of non-rigid registration in the 4DMatch/4DLoMatch benchmark. The top two lines are from 4DMatch, while the bottom three are from 4DLoMatch. The blue and yellow colors denote the source and target point cloud, respectively. The green and red lines indicate whether the predicted deformable flow from the source points is accepted by the threshold. The deformable registration is built by GraphSCNet [25]. Zoom in for details.